How many possibilities are there to select a team of 3 members from 9 players?
Answer by edjones(8007)

nCr=combination of n things taken r at a time=n!/(n-c)!*c!
9!/(6!*3!)
=9*8*7*6!/6!*6
=(9*8*7)/6=84
.
Ed